OpenStudy (anonymous):

larry wants to buy some carpeting for his living room. The length of the room is 2 times the width and the total area of the room is 18 square meters. What is the length of the living room?

OpenStudy (chris):

They want you to turn the word problem into a set of equations - you will need two equations since you have two unknowns (the length and width of the living room).

OpenStudy (chris):

The first equation relates length to width: 1. l = 2w (the length is twice the width) The second equation tells you the area, which we know is l*w for any rectangle. 2. lw = 18

OpenStudy (chris):

Do you know how to solve it from there?

OpenStudy (chris):

One way is to "substitute" the first equation into the second equation. So since the first equation says that l = 2w, we replace "l" in the second equation with "2w": 3. 2w*w = 18 4. 2w^2 = 18 5. w^2 = 18/2 6. w^2 = 9 7 w = sqrt(9) 8. w = 3. So now you know the width of the room is "3" meters. Now substitute that back into either of the first two equations and solve for l, the length.

OpenStudy (chris):

1. l = 2w 2. l = 2*3 3. l = 6 The length of the room is 6

OpenStudy (chris):

I'll leave it to you to check the work - given l = 6 and w = 3, does the word problem make sense when you plug in those values (do both equations work?).

OpenStudy (chris):

Let me know if you're confused.